OpenStudy (anonymous):

Factor the polynomial f(x). Then solve the equation f(x)=0. f(x)=x^3-3x^2-10x+24

hero (hero):

By Rational Root Theorem, the likely zeroes are, 1,2,3,4,6,8,12,24. After a bit of trial and error, we see that x = 2, is a zero, therefore x - 2 = 0 and x - 2 is a factor of f(x).

hero (hero):

Next, Take f(x) and split it in accordance with the known factor x - 2, then factor the remaining quadratic as follows: x^3 - 3x^2 - 10x + 24 = x^3 - 2x^2 - x^2 + 2x - 12x + 24 = x^2(x - 2) - x(x - 2) - 12(x - 2) = (x - 2)(x^2 - x - 12) <---remaining quadratic = (x - 2)(x - 4)(x + 3)

hero (hero):

So f(x) = (x - 2)(x - 4)(x + 3) Now set f(x) = 0 0 = (x - 2)(x - 4)(x + 3) By Zero Product Property: x - 2 = 0 x - 4 = 0 x + 3 = 0 You should be able to find x from there.

hero (hero):

Any questions?

OpenStudy (anonymous):

Thank you

hero (hero):

You're welcome